OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (-2,-3) and has a slope of 1/2

OpenStudy (anonymous):

y-y1=m(x-x1) x1 = -2 y1 = -3 m = 1/2

OpenStudy (anonymous):

y-(-3)=0.5(x-(-2)) y+3=0.5(x+2) y+3=0.5x+1 y=0.5x-2